OpenStudy (anonymous):

The function for the cost of materials to make a hat is f(x) = one half x + 1, where x is the number of hats. The function for the selling price of those hats is g(f(x)), where g(x) = x + 2. Find the selling price of 10 hats. 6 8 10 12

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

g(f(x))=x+2 =[1/2x+1] +2 =[1/2(10)+1] +2 =6+2=8

OpenStudy (anonymous):

okay what else